OpenStudy (mendicant_bias):

(Orthogonal Functions)(Complex Functions) I'm having trouble properly determining the norm of a given function; I'm getting zero, and that obviously can't be right. Workings posted below shortly.

OpenStudy (mendicant_bias):

\[\text{Determine the norm in set (a).} \ \ \ f(x)=e^{inx}.\]

OpenStudy (mendicant_bias):

\[||f(x)||^2=\int\limits_{-\pi}^{\pi}f(x)^2dx=\int\limits_{-\pi}^{\pi}\bigg(e^{inx}\cdot e^{inx}\bigg)dx=\int\limits_{-\pi}^{\pi}\bigg(e^{2inx}\bigg)dx\]

OpenStudy (mendicant_bias):

\[=\frac{e^{2inx}}{2 i n}\bigg]_{-\pi}^{\pi}=\frac{e^{i2\pi n}-e^{-i2\pi n}}{2 i n}\]

OpenStudy (mendicant_bias):

If you expand that using Euler's formula, every term in the numerator either cancels out or becomes zero upon evaluating them as trig functions, so I clearly must be doing something wrong. Where is my mistake? @wio

OpenStudy (mendicant_bias):

@perl

OpenStudy (perl):

you dont need to expand it

OpenStudy (mendicant_bias):

But it's not to pi i n, it's to 2 pi i n, unless you can just separate them and leave e^2.

OpenStudy (perl):

you can treat the two cases, when n is odd and when n is even. in both cases

OpenStudy (perl):

e^(2 pi * i * n) = 1, for any n

OpenStudy (perl):

e^(- 2 pi * i * n) = 1, for any n

OpenStudy (mendicant_bias):

Alright, so then you just have 1-1=0 in the numerator?

OpenStudy (perl):

yes

OpenStudy (mendicant_bias):

But then your norm would be zero, there's no way that can be correct. The norm isn't supposed to be zero, right? That doesn't make any sense.

OpenStudy (perl):

e^(2 pi * i * n) , this is the point (1,0) when n=0, that rotates 360 degrees every time you increment n by 1. e^( pi * i * n) , this is the point (1,0) when n=0, that rotates 180 degrees every time you increment n by 1.

OpenStudy (mendicant_bias):

So we're saying a function is orthogonal to itself.

OpenStudy (perl):

yes, `an orthogonal function` is orthogonal to itself

OpenStudy (perl):

not just any function is orthogonal to itself

OpenStudy (mendicant_bias):

What you are saying is making little to no sense to me. What is an "orthogonal function"? If something is orthogonal, it happens to be orthogonal to something else, a set might behave orthogonally, but it's not like you have these functions and they can just be called "orthogonal". They're orthogonal under certain conditions, and it makes absolutely no sense to me, with the extension of the idea of orthogonality from vectors to functions, that a function can be orthogonal to itself. I've never heard of a vector being orthogonal to itself, and I don't see neither how a function can just be called "orthogonal" (unless it's very explicitly used to reference a set of functions being orthogonal to one another), nor how a function can be orthogonal to itself.

OpenStudy (mendicant_bias):

I've just never seen or heard, to my knowledge, of a function being orthogonal to itself, so I'm having a hard time believing this. And also the way you used orthogonal threw me off a lot. I don't understand.

OpenStudy (perl):

$$\Large { \text{In mathematics, two functions f and g}\\ \text { are called orthogonal} \text{ if their inner product }\\{ \langle f,g\rangle} \text{ is zero for } f \neq g \text{ where } \\ \langle f,g\rangle=\int_\mathbb{C} \overline {f(x)} . g(x) }$$ http://en.wikipedia.org/wiki/Orthogonal_functions

OpenStudy (mendicant_bias):

Yep. I just haven't seen until now a function that has a norm of zero. To my understanding that would be like having a vector of zero length. It's trivial, it's meaningless, it's supposed to have some value, because you're supposed to use that norm to make that a unit vector, or in this instance, to make something orthonormal. No disagreement on the definition.

OpenStudy (mendicant_bias):

Feeling 99% sure that a norm of zero absolutely should not be the answer here, in any case. This just doesn't make any sense. Unless you're dealing with something that's already like, the zero vector, taking the norm of it, as a vector or otherwise, shouldn't give you zero. The length of said vector is not zero, and neither should generally be the inner product of a function with itself evaluated over some bounds.

OpenStudy (mendicant_bias):

"From the linearity property it is derived that x = 0 implies \langle x,x \rangle = 0, while from the positive-definiteness axiom we obtain the converse, \langle x,x \rangle = 0 implies x = 0. Combining these two, we have the property that \langle x,x \rangle = 0 if and only if x = 0." That last sentence there.

OpenStudy (perl):

page 8

OpenStudy (mendicant_bias):

Those are two different functions, v_j =/= v_k. That's not taking the inner product of a function with itself.

OpenStudy (mendicant_bias):

Yep, in agreement there entirely.

OpenStudy (perl):

length and orthogonality are different concepts

OpenStudy (mendicant_bias):

Either way, why did it pretty overwhelmingly straightforwardly state as a property of the inner product that *unless your function itself is zero*, its norm should not be zero. I mean, that's literally what it says. Is that or is that not true? Does it say something else? I can't think of any examples where the norm of a nonzero functionwould be zero, and I can't find any.

OpenStudy (mendicant_bias):

Like, 100%, straightforward, simple yes or no, can the norm of a nonzero function be zero? Absolutely not is my current understanding, and-again-other examples would be fantastic, I am literally just failing to imagine any example, any at all, where the norm of a nonzero function is zero. Somebody has to have an example. I feel like I'm taking crazy pills.

OpenStudy (perl):

lets solve < f(x), f(x)> ^(1/2)

OpenStudy (mendicant_bias):

Can I get just a super straightforward, word answer to my question about the comment on the Inner Product wikipedia page.

OpenStudy (mendicant_bias):

It gets super infuriating in that people on here have a habit of throwing a ton of math uo-and that's fantastic, but I just need an absolutely firm, clear answer, is that just wrong? Am I misreading it? Is it wrong? What's the issue here? Is that entry wrong, or am I misunderstanding it?

OpenStudy (mendicant_bias):

Can you please answer my question regarding the Wiki entry so we can clear that up and move on.

OpenStudy (perl):

ok this is what i wanted $$ \Large {{ || f(x)|| = } { \langle f(x), f(x) \rangle }^{\frac{1}{2}} } $$

OpenStudy (mendicant_bias):

Alright, cool, can you answer my question on the Wiki entry, it's a really simple, worded, "yes" or "no" answer.

OpenStudy (perl):

one sec

OpenStudy (perl):

I don't think so, but i am checking

OpenStudy (mendicant_bias):

I don't understand that why, for the norm, you are integrating from zero to one, my book just has a to be, not ever zero to one.

OpenStudy (mendicant_bias):

*a to b as generic values of choice

OpenStudy (perl):

ok i was thinking of a special case of orthogonal functions, fourier series

OpenStudy (mendicant_bias):

Yeah, no worries, I'm not dealing with that, though, I'm just saying, as a general statement: If you have a nonzero function and you find its norm, the norm should not be zero. This function that I started with, to my understanding, is a nonzero function, and I do not understand how its norm is yielding zero, then, and am assuming an algebraic error. The only way that the function that I started with, that we've been having this whole rambling discussion over, can have a norm of zero, is if it itself is zero, according to the Wikipedia page section under elementary properties. I do not understand how apparently e^{inx}=0.

OpenStudy (perl):

im reading this http://tutorial.math.lamar.edu/Classes/DE/PeriodicOrthogonal.aspx it seems that we are looking at $$ \Large {\{f_n(x)}\} $$

OpenStudy (perl):

scroll down to the definition

OpenStudy (perl):

so we want to show that $$\Large {\{ e^{\pi i n }} \} ,~ n=0,-1,1,-2,2,-3,3,...$$ is mutually orthogonal on (-pi, pi )

OpenStudy (mendicant_bias):

All we're trying to do is find the norm, that's all.

OpenStudy (perl):

ok let me check on wolfram, be right back

OpenStudy (mendicant_bias):

Looks like it's nonzero on Wolfram Alpha, look at that. One hour down the drain. One hour of both of our time. I should've just gone to W|A in the first place.

OpenStudy (perl):

can you give me link

OpenStudy (perl):

yes because you have to take the complex conjugate when you do dot product

OpenStudy (mendicant_bias):

It's 6 AM.

OpenStudy (mendicant_bias):

I'm going to bed, I have literally no idea whatsoever why it took nearly two hours to point out that I wasn't using the proper definition, but thanks, I guess. At least that was figured out. After two hours.

OpenStudy (mendicant_bias):

(Oh, sorry, correction, I asked two hours ago, we started talking over an hour ago. My bad. It just took two hours for that to get pointed out, but you had only been helping me out for an hour. Thanks, perl. Have a good day.)

OpenStudy (perl):

$$ \Large {{ \langle f(x), f(x) \rangle } = \int\limits_{-\pi}^{\pi}f(x) ~f(x) ^{*}dx \\=\int\limits_{-\pi}^{\pi}\bigg( e^{inx}\cdot \overline {e^{inx}} ~\bigg ) dx =\int\limits_{-\pi}^{\pi}\bigg( e^{inx}\cdot {e^{-inx}} ~\bigg ) dx\\ =\int\limits_{-\pi}^{\pi}\bigg( e^{inx-inx} ~\bigg ) dx\\ =\int\limits_{-\pi}^{\pi}\bigg( e^{0} ~\bigg ) dx\\ =\int\limits_{-\pi}^{\pi}\bigg( 1 ~\bigg ) dx\\ =\left[x \right]_{\hbox{-}\pi}^{\pi} = \pi - (-\pi) = 2\pi } $$

OpenStudy (perl):

the norm is the square root of this

OpenStudy (perl):

so all these functions have length sqrt(2 Pi ) and these functions are mutually orthogonal , when n does not equal to m . so dont use wolfram here, it would be inappropriate. unless you use the proper definition using the complex conjugate

OpenStudy (perl):

http://www.wolframalpha.com/input/?i=evaluate+integral+e^%28nix%29*e^%28conjugate%28i%29*nx%29%2C+x+%3D+-pi..+pi Note that if you conjugated the entire expression on the second term you would get a strange integral. wolfram does not know you are only interested in real x values on eals [0,1] and n is an integer. strange integral (note the conjugation operator ^* ): http://www.wolframalpha.com/input/?i=evaluate+integral+e^%28nix%29*%28e^%28i*n*x%29%29^*%2C+x+%3D+-pi..+pi

OpenStudy (mendicant_bias):

^I'm way more inclined to believe your result than the W|A output, do you have any idea how they got that? Also, thank you.